OpenStudy (anonymous):

Lenny started making contributions to a Traditional IRA when he got his first job, at age 24. His contributions averaged $2,400 annually. Lenny was in the 25% tax bracket during his working years, when he continued to make these annual contributions. The average annual rate of return on the account was 6.5%. Upon retirement at age 67, Lenny stopped working and making IRA contributions. Instead, he started living on withdrawals from the retirement account. At this point, Lenny dropped into the 15% tax bracket. Factoring in taxes, what is the effective value of Lenny’s Traditional IRA at retirem

OpenStudy (anonymous):

Assume annual compounding

OpenStudy (anonymous):

@ganeshie8 @*Scotto0715*

ganeshie8 (ganeshie8):

C = 2400 r = 0.065 t = 67-24

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

use the annuity formula and find out the accumulated amount

ganeshie8 (ganeshie8):

Future Value = \(\large C \bullet \frac{(1+i)^{nt}-1}{i}\)

ganeshie8 (ganeshie8):

since this is annual compounding, here, i = r/n = 0.0625/1 = 0.0625

ganeshie8 (ganeshie8):

plugin the numbers and evaluate

ganeshie8 (ganeshie8):

Future Value = \(\large 2400\bullet \dfrac{(1+0.065)^{67-24}-1}{0.065}\)

ganeshie8 (ganeshie8):

you need to subtract taxes on above amount

OpenStudy (anonymous):

answer is $439,321.62